# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v2

@@ Abstract
The gravitational waveform emitted by a particle moving in a Schwarzschild geometry can be computed, in perturbation theory, as a double series expansion in the limit where distances are large with respect to the size of the horizon, $z=2M/r\ll 1$ (Post-Minkowskian approximation), but small with respect to the gravitational wavelength, $y=2{\rm i}\omega r\ll 1$ (Post-Newtonian approximation). While in the case of bounded systems the two expansions are linked, for scattering processes they describe complementary regions of spacetime. In this paper, we derive all order formulae for those two expansions of the waveform, where the entire dependence on the $y$ or $z$ variables is resummed under the assumption $\omega M\ll 1$. The results are based on a novel hypergeometric representation of the confluent Heun functions and provide a new derivation of the recently discovered Heun connection formulae. Similar results are found for the non-confluent Heun case. Finally, in the case of circular orbits we compute our formulae at order 30PN and show excellent agreement against solutions obtained via numerical integrations and the MST method.
I Introduction
@@
Recent advances in the sensitivity of experimental apparatus for detecting gravitational waves (GWs) necessitate increasingly precise theoretical computations. The next generation of space-based experiments, such as the Laser Interferometer Space Antenna (LISA), promises to revolutionize the GW detection by significantly enhancing sensitivity and exploring entirely new wavelength ranges compared to ground-based interferometers [1]. LISA will detect long-wavelength GWs originating from sources like supermassive black hole mergers, extreme mass-ratio inspirals (EMRIs), galactic binaries, and the Stochastic Gravitational Wave Background—sources largely inaccessible to ground-based observatories.
@@
The gravitational signals from these sources are diverse. For instance, GWs from supermassive black hole mergers will produce strong signals with large signal-to-noise ratios, while those from EMRIs, though weaker, will carry exceptionally detailed information about the gravitational sources. Indeed, EMRIs waveforms, spanning thousands of cycles during their inspiral phase, provide a unique opportunity to probe fundamental questions about the nature of gravitational sources, including their compactness, the presence or absence of horizons, light rings, echoes [2], and the validity of Einstein’s theory of gravity and its possible extensions [3, 4].
@@
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion [5, 6, 7] and perturbation theory [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in [20]. The expansion is known to be converging slowly [21, 9] due to the presence of logarithmic terms in the velocity, see also [22, 23, 24, 25]. In [26, 27, 28, 29, 30, 31, 32, 33, 34] novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a $SU(2)^{2}$ quiver gauge theory. Following this path the infinite tower of logarithmic terms (tails and tails of tails) have been resummed into exponentials.
@@
More recently, multipole [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and perturbation theory [49] methods have been applied to the study of scattering processes (open trajectories) in gravity [50, 51], a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see [52] for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
@@
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits $x\to 0$ and $y$ or $z$ finite lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN approximations respectively. The parameters of the hypergeometric function and the coefficients of the polynomials are all expressed in terms of a characteristic function $a(\ell)$ that, in the MST [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] formulation of the confluent Heun solution, is the ”renormalized angular momentum”, while in the description of this paper [26, 27, 28, 29, 30, 31, 32, 33, 34, 49], it is the Seiberg-Witten (SW) period of the ${\cal N}=2$ supersymmetric gauge theory. In this framework the gauge couplings of the quiver gauge theory are connected to the PM and PN expansion parameters as $q_{1}=z$ and $q_{2}=y$. The same techniques apply to D-branes, topological stars and other gravitational backgrounds governed by an equation of the Heun type. See [53, 54] for a recent application to the case of topological stars.
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. These results provide a first principle proof of the Heun connection formula derived in [26, 27, 28, 29] from the crossing symmetry of the conformal blocks of a conformal field theory (CFT) . Our results provide an algorithm which is very efficient when implemented in a Mathematica code and that will allow us to compute the gravitational waveform in the case of circular orbits up to the 30th PN order. This latter result will be tested against a numerical solution and the 22nd PN order computation available in literature [20]. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section II we derive the PM and PN expansion of the confluent Heun function. In Section III we apply the results of Section II to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section IV we test our results against those obtained via the MST method and numerical integration. In Section V we extend the results to include the case of the non-confluent Heun function. Section VI contains our conclusions.
II The confluent Heun equation
@@
$g^{\prime\prime}(t)+g^{\prime}(t)\left(\zeta+\frac{\gamma}{t}+\frac{\delta}{t-1}\right)+\frac{g(t)(\beta t-q)}{(t-1)t}=0$ (1)
@@
${\rm HeunC}[q,\beta,\gamma,\delta,\zeta,t]$ (2)
$t^{1-\gamma}\text{HeunC}[(1-\gamma)(\zeta-\delta)+q,\beta+(1-\gamma)\zeta,2-\gamma,\delta,\zeta,t]$
with ${\rm HeunC}$ the confluent Heun function. It is convenient to bring the irregular singularity to the origin, by taking $t=1/z$, and introduce the parameters $u$, $x$, $\kappa_{i}$ (with $i=1,2,3$) for the confluent Heun labels
$\beta$ $=$ $-x(\kappa_{1}+\kappa_{3}),\quad\delta=\kappa_{2}+\kappa_{3},\quad\gamma=1-\kappa_{2}+\kappa_{3}$
$\zeta$ $=$ $-x\qquad,\qquad\quad q=u-\kappa_{3}^{2}-x(\kappa_{1}+\kappa_{3})$ (3)
In these variables, the confluent Heun equation (CHE) takes the form
$G^{\prime\prime}(z){+}G^{\prime}(z)\left(\frac{x}{z^{2}}{+}\frac{2\kappa_{1}{+}1}{z}{+}\frac{\kappa_{2}{+}\kappa_{3}}{z{-}1}\right){+}G(z)\frac{\left(u{-}\kappa_{1}^{2}{+}\left(\kappa_{1}{+}\kappa_{2}\right)\left(\kappa_{1}{+}\kappa_{3}\right)z\right)}{(z{-}1)z^{2}}=0$ (4)
with
$G(z)=z^{-\kappa_{1}-\kappa_{3}}g({\textstyle\frac{1}{z}})$ (5)
There are two limits in which (4) admits a simple solution. They correspond to sending $x\to 0$ keeping finite either $z$ or the ratio
$y={x\over z}$ (6)
In these limits (4) reduces to a hypergeometric equation with solutions
$G_{\alpha}(z)\underset{x,z\to 0\atop x/z\leavevmode\nobreak\ {\rm finite}}{\approx}H^{0}_{\alpha}({\textstyle\frac{x}{z}})\qquad,\qquad G_{\alpha}(z)\underset{x\to 0\atop z\leavevmode\nobreak\ {\rm finite}}{\approx}H^{1}_{\alpha}(z)$ (7)
where $\alpha=\pm$ labels the two independent solutions,
$H^{0}_{\alpha}(y)$ $=$ $y^{-\alpha a+\kappa_{1}}{}_{1}F_{1}(\kappa_{1}{-}\alpha a,1{-}2\alpha a;y)$
$H^{1}_{\alpha}(z)$ $=$ $z^{\alpha a-\kappa_{1}}{}_{2}F_{1}\left(\alpha a{+}\kappa_{2},\alpha a{+}\kappa_{3};1{+}2\alpha a;z\right)$ (8)
and
$a=\sqrt{u}+O(x)$ (9)

### v3

@@ Abstract
We derive formulae that resum, at a given order in the limit of small frequency, the infinite series of Post-Minkowkian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in Schwarzschild geometry. Specifying to the case of circular orbits, we compute the wave form and the energy flux to order 30PN, and compare it against the available results in the literature. The results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a mathematical proof of the Heun connection formula.
1 Introduction
@@
Recent advances in the sensitivity of experimental apparatus for detecting gravitational waves (GWs) necessitate increasingly precise theoretical computations. The next generation of space-based experiments, such as the Laser Interferometer Space Antenna (LISA), promises to revolutionize the GW detection by significantly enhancing sensitivity and exploring entirely new wavelength ranges compared to ground-based interferometers LISA:2022kgy . LISA will detect long-wavelength GWs originating from sources like supermassive black hole mergers, extreme mass-ratio inspirals (EMRIs), galactic binaries, and the Stochastic Gravitational Wave Background—sources largely inaccessible to ground-based observatories.
@@
The gravitational signals from these sources are diverse. For instance, GWs from supermassive black hole mergers will produce strong signals with large signal-to-noise ratios, while those from EMRIs, though weaker, will carry exceptionally detailed information about the gravitational sources. Indeed, EMRIs waveforms, spanning thousands of cycles during their inspiral phase, provide a unique opportunity to probe fundamental questions about the nature of gravitational sources, including their compactness, the presence or absence of horizons, light rings, echoes Cardoso:2019rvt , and the validity of Einstein’s theory of gravity and its possible extensions Gair:2012nm ; Yunes:2013dva .
@@
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion Thorne:1980ru ; Damour:1982wm ; Blanchet:2013haa and perturbation theory Teukolsky:1973ha ; Poisson:1993vp ; Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk . The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
@@
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
@@
More recently, multipole Bini:2023fiz ; Brandhuber:2023hhy ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Herderschee:2023fxh ; Georgoudis:2023eke ; Georgoudis:2023lgf ; Georgoudis:2023ozp ; Elkhidir:2023dco ; Bini:2024ijq ; Bini:2024rsy ; Georgoudis:2024pdz ; Alessio:2024onn ; Brunello:2024ibk and perturbation theory Fucito:2024wlg methods have been applied to the study of scattering processes (open trajectories) in gravity Kovacs:1977uw ; Kovacs:1978eu , a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see DiVecchia:2023frv for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
@@
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits where $x$ is taking to be small keeping finite $y$ or $z$ lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN expansions respectively.
We observe that, the hypergeometric representation of the confluent Heun function we derived here is similar in spirit to the well studied MST solution Mano:1996vt . The crucial difference between the two, is that while the MST ansatz involves an infinite sum of hypergeometric functions, our solution is written in terms of a single hypergeometric function and its first derivative.
This leads to a mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The whole non-triviality of the Heun equation (and its confluences) is codified in a single function, $a(\ell)$, the Seiberg-Witten (SW) period of the underlying ${\cal N}=2$ supersymmetric gauge theory (see also Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe ). This function is related to the so called "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk or "anomalous dimension" in the effective field theory context Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resuming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to the 30th PN order [^2] The ancillary mathematica files coming with this paper should help the reader to check our computations.. This result will be tested against the 22nd PN order computation available in literature Fujita:2012cm , and results based on MST and numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit . The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. In Section 5 we extend the results to include the case of the non-confluent Heun function. Section 6 contains our conclusions.
2 The confluent Heun equation
@@
$g^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right)g^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}g(t)=0$ (1)
@@
${\rm HeunC}[\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},\lambda,t]$ (2)
$t^{1-\beta_{0}}\text{HeunC}[\alpha_{0}+(1-\beta_{0})(\lambda-\beta_{1}),\alpha_{1}+\lambda(1-\beta_{0}),2-\beta_{0},\beta_{1},\lambda,t]$
with ${\rm HeunC}$ the confluent Heun function.
2.1 Confluent Heun-Gauge dictionary
The confluent Heun equation can be viewed as the quantum version of the Seiberg-Witten (SW) curve governing the dynamics of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive flavours ($m_{i}$)transforming in the fundamental representation of the gauge group (with gauge coupling $x$) Bianchi:2021mft ; Bianchi:2021xpr living on a curved Nekrasov-Shatashvilli background Nekrasov:2009rc . This suggests a a change of variables in the Heun equation (1) in terms of the parameters of the SW quantum curve and its Coulomb branch parameter $u$
$\alpha_{0}$ $=$ $u-{\textstyle\frac{1}{4}}+m_{1}-m_{1}^{2}+(m_{1}+m_{3}-1)x\qquad,\qquad\alpha_{1}=\left(m_{1}+m_{3}-1\right)x$
$\beta_{0}$ $=$ $1-m_{1}+m_{2}\qquad,\qquad\beta_{1}=1-m_{1}-m_{2}\qquad,\qquad\lambda=-x$ (3)
which will simplify the computations which we are going to perform later in the paper. Defining
$G_{1}(z)=z^{-{\alpha_{1}\over\lambda}}g({\textstyle\frac{1}{z}})$ (4)
the Heun equation (1) can be put into the form [^3] Alternatively, equation (8) can be written in the normal Schrödinger like form $\Psi^{\prime\prime}(z)+Q_{\rm gauge}(z)\Psi=0\qquad,\qquad G_{1}(z)=e^{{\frac{x}{2z}}}z^{m_{3}-1}(1-z)^{m_{1}+m_{2}-1\over 2}\Psi(z)$ (5) with $Q_{\rm gauge}={-}\frac{x^{2}}{4z^{4}}{+}\frac{xm_{3}}{z^{3}}{+}\frac{1{-}\left(m_{1}{-}m_{2}\right)^{2}}{4(z{-}1)z}+\frac{1{-}\left(m_{1}{+}m_{2}\right){}^{2}}{4(z-1)^{2}z}+\frac{u{-}{\textstyle\frac{1}{4}}{+}{\textstyle\frac{x}{2}}\left(m_{1}{+}m_{2}{-}1\right)}{(z{-}1)z^{2}}$ (6) Notice that also the conformal Ward identity of Liouville theory, the so called BPZ equation, is a confluent Heun equation. The normal form of such equation matches again with (5) and (6)after the identifications $m_{1}=p_{0}-k_{0}\qquad,\qquad m_{2}=-p_{0}-k_{0}\qquad,\qquad m_{3}=c\qquad,\qquad x=-x_{\rm CFT}$ (7) with $k_{0}$, $p_{0}$, $c$ and $x_{\rm CFT}$ the CFT variables in Consoli:2022eey ; Fucito:2023afe .
$G_{1}^{\prime\prime}(z)+G_{1}^{\prime}(z)\left(\frac{1{-}m_{1}{-}m_{2}}{z{-}1}{-}\frac{2\left(m_{3}{-}1\right)}{z}{+}\frac{x}{z^{2}}\right)$ (8)
$\qquad{+}G_{1}(z)\left(\frac{{\textstyle\frac{1}{4}}{-}u{+}m_{3}(m_{3}{-}1)}{z^{2}}{+}\frac{u+\frac{3}{4}-m_{1}-m_{2}-m_{3}+m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}}{(z{-}1)z}\right)=0$
where now the irregular point is located at the origin. Alternatively, one writes
$G_{0}(z)=G_{1}({\textstyle\frac{x}{z}})$ (9)
and setting $y=x/z$ the confluent Heun equation (8) becomes
$G_{0}^{\prime\prime}(y){+}G_{0}^{\prime}(y)\left(\frac{1{-}m_{1}{-}m_{2}}{y{-}x}{+}\frac{m_{1}{+}m_{2}{+}2m_{3}{-}1}{y}{-}1\right)$ (10)
$\qquad\qquad{+}G_{0}(y)\left(\frac{u{+}{\textstyle\frac{3}{4}}{+}\delta}{(x-y)y}{+}\frac{1{+}\delta{+}m_{3}(m_{3}{-}1)}{y^{2}}\right)=0$
where $\delta=m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}-m_{1}-m_{2}-m_{3}$. We will use both descriptions of the confluent Heun equation, and we will refer to them as CHE. In particular, we will consider the expansions of the two equations in the limit where $x$ is taken small but keeping finite either $z$ or $y=x/z$. The two expansions will correspond in the gravity context to the Post-Minkowskian and Post-Newtonian approximations, so we will refer to $G_{0}$ and $G_{1}$ as the PM and PN representations of the confluent Heun function respectively. In particular, setting $x\to 0$, (10) and (8) reduce to hypergeometric equations with solutions
${\rm PM}:$ $\qquad G^{0}_{\alpha}(y)\underset{x\to 0\atop y~{}{\rm finite}}{\approx}H^{0}_{\alpha}(y)$
${\rm PN}:$ $\qquad G^{1}_{\alpha}(z)\underset{x\to 0\atop z~{}{\rm finite}}{\approx}H^{1}_{\alpha}(z)$ (11)
with $\alpha=\pm$ labeling the two independent solutions,
$H^{0}_{\alpha}(y)$ $=$ $y^{\frac{1}{2}-\alpha a-m_{3}}\,_{1}F_{1}\left({\textstyle\frac{1}{2}}-\alpha a-m_{3};1-2\alpha a;y\right)$
$H^{1}_{\alpha}(z)$ $=$ $z^{-\frac{1}{2}+\alpha a+m_{3}}\,_{2}F_{1}\left(\frac{1}{2}+\alpha a-m_{1},\frac{1}{2}+\alpha a-m_{2};1+2\alpha a;z\right)$ (12)
and $a\underset{x\to 0}{\approx}\sqrt{u}$. On the other hand, at order $k$ in $x$, we will look for solutions of the form
${\rm PM}_{k}:$ $\qquad G^{0}_{\alpha}(y)=y^{\frac{1}{2}-\alpha a-m_{3}}\sum_{i=0}^{k}\sum_{j=0}^{\infty}c^{PM}_{ij}(\alpha a)x^{i}y^{j-i}$
${\rm PN}_{k}:$ $\qquad G^{1}_{\alpha}(z)=z^{-\frac{1}{2}+\alpha a+m_{3}}\sum_{j=0}^{k}\sum_{i=0}^{\infty}c^{PN}_{ij}(\alpha a)z^{i-j}x^{j}$ (13)
where $a(u)$ is a function of $u$. Writing
$u=a^{2}+\sum_{i=1}^{k}u_{i}(a)x^{i}$ (14)
one can determine the coefficients $c^{PM}_{ij}(\alpha a)$, $c^{PN}_{ij}(\alpha a)$ and $u_{i}(a)$ by solving the CHE equations order by order in $x$. We notice that $a$ codify the monodromy of the solutions around the irregular point $z=0$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$
${\rm Near~{}zone}:\qquad G^{1}_{\alpha}(z)=g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ (15)
The main result of this Section is that, the expansions PMk and PNk in (2.1) can be exactly resumed and written in terms of a single hypergeometric function. This will provide us with a first principle proof of the connection formulae for the confluent Heun functions, where the whole non-triviality of the Heun equation is codified into a single function $a(u)$.